# PRIMA.CPP: Speeding Up 70B-Scale LLM Inference on Low-Resource Everyday Home Clusters

### v1

@@ A.3 Layer-to-Device Assignment: From Latency Analysis to Vectorized Model
To apply constraints to the subset of ${\bm{w}}$ and ${\bm{n}}$ corresponding to $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}$, we define diagonal matricies ${\bm{P}}_{w}=\operatorname{diag}(-{\bm{I}}_{\mathcal{M}_{1}},-{\bm{I}}_{\mathcal{M}_{2}},-{\bm{I}}_{\mathcal{M}_{3}},{\bm{P}}_{\mathcal{M}_{4}}^{1},{\bm{P}}_{\mathcal{M}_{4}}^{2},{\bm{P}}_{\mathcal{M}_{4}}^{3})$, ${\bm{P}}_{n}=\operatorname{diag}(\mathbf{0}_{\mathcal{M}_{1}},\mathbf{0}_{\mathcal{M}_{2}},{\bm{I}}_{\mathcal{M}_{3}},\mathbf{0}_{\mathcal{M}_{4}}^{1},\mathbf{0}_{\mathcal{M}_{4}}^{2},-\mathbf{P}_{\mathcal{M}_{4}}^{3})$, where ${\bm{I}}_{\mathcal{M}_{1}},{\bm{I}}_{\mathcal{M}_{2}},{\bm{I}}_{\mathcal{M}_{3}}$ are identity matrices and $\mathbf{0}_{\mathcal{M}_{1}},\mathbf{0}_{\mathcal{M}_{2}},\mathbf{0}_{\mathcal{M}_{3}}$ are zero matrices corresponding to the subsets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$, and ${\bm{P}}_{\mathcal{M}_{4}}^{1},{\bm{P}}_{\mathcal{M}_{4}}^{2},{\bm{P}}_{\mathcal{M}_{4}}^{3}$ are diagonal binary matricies (i.e., selection matricies) corresponding to the three constraints (31-33) within the subset $\mathcal{M}_{4}$. To construct ${\bm{P}}_{\mathcal{M}_{4}}^{1},{\bm{P}}_{\mathcal{M}_{4}}^{2},{\bm{P}}_{\mathcal{M}_{4}}^{3}$, we define a binary vector ${\bm{p}}_{\text{macOS}}$, where a value of 1 indicates that the current device is running on macOS and a value of 0 indicates otherwise. The number of elements in ${\bm{p}}_{\text{macOS}}$ matches the number of devices in the set $\mathcal{M}_{4}$. Similarly, we define binary vectors ${\bm{p}}_{\text{Linux}}$, ${\bm{p}}_{\text{Android}}$, ${\bm{p}}_{\text{metal}}$. Thus, we have ${\bm{P}}_{\mathcal{M}_{4}}^{1}=\operatorname{diag}({\bm{p}}_{\text{macOS}}\odot(1-{\bm{p}}_{\text{metal}}))$, ${\bm{P}}_{\mathcal{M}_{4}}^{2}=\operatorname{diag}({\bm{p}}_{\text{macOS}}\odot{\bm{p}}_{\text{metal}})$, ${\bm{P}}_{\mathcal{M}_{4}}^{3}={\bm{p}}_{\text{Linux}}+{\bm{p}}_{\text{Android}}$.
To handle constraints (35-36), we define ${\bm{P}}_{n}^{\text{gpu}}$ as a similar diagonal binary matrix, with elements set to one for devices with CUDA or Metal support. Specifically, we let ${\bm{P}}_{n}^{\text{gpu}}={\bm{P}}_{n}^{\text{cuda}}+{\bm{P}}_{n}^{\text{metal}}$, where ${\bm{P}}_{n}^{\text{cuda}}=\operatorname{diag}(\mathbf{0}_{\mathcal{M}_{1}},\mathbf{0}_{\mathcal{M}_{2}},{\bm{P}}_{\mathcal{M}_{3}}^{\text{cuda}},{\bm{P}}_{\mathcal{M}_{4}}^{\text{cuda}})$ and ${\bm{P}}_{n}^{\text{metal}}=\operatorname{diag}(\mathbf{0}_{\mathcal{M}_{1}},{\bm{I}}_{\mathcal{M}_{2}},\mathbf{0}_{\mathcal{M}_{3}},{\bm{P}}_{\mathcal{M}_{4}}^{\text{metal}})$.
@@
and the VRAM/shared memory upper bound be ${\bm{z}}^{\text{gpu}}=[z_{1}^{\text{gpu}},\cdots,z_{M}^{\text{gpu}}]$, where
$z_{m}^{\text{gpu}}=\frac{1}{Lb^{\prime}}\cdot\begin{cases}0,&\text{if}\ \ \mathbb{I}_{\text{cuda}}=0\,\,\text{and}\,\,\mathbb{I}_{\text{metal}}=0,\\ d_{m,\text{cuda}}^{\text{avail}}-c^{\text{gpu}},&\text{if}\ \ \mathbb{I}_{\text{cuda}}=1,\\ d_{m,\text{metal}}^{\text{avail}}-c^{\text{gpu}},&\text{if}\ \ \mathbb{I}_{\text{metal}}=1\,\,\text{and}\,\,m\neq 1,\\ d_{m,\text{metal}}^{\text{avail}}-c^{\text{gpu}}-d_{o},&\text{if}\ \ \mathbb{I}_{\text{metal}}=1\,\,\text{and}\,\,m=1.\end{cases}$
The problem model can then be reformated as:
@@
$b,b_{i},b_{o}$ Bytes of weight data for each layer, input, and output.
$n_{kv}$ Number of tokens stored in key-value cache.
@@
$s_{\text{threshold}}^{\text{disk}}$ A threshold for disk speed. If the disk speed is below this threshold, it is considered too slow.
@@
${\bm{P}}_{w},{\bm{P}}_{n}$ Constraint coefficients for $w_{m}$ and $n_{m}$, should be diagonal matrices.
${\bm{P}}_{n}^{\text{gpu}}$ A diagonal binary matrix that indicates whether a device uses a GPU.
@@
A.4 Run Prima.cpp with More Hot Models: Llama, Qwen, QwQ and DeepSeek
Figure 7 provides a clear comparison, showing that for models larger than 30B, our prima.cpp always has the lowest token latency and TTFT. We also support Qwen-2.5, QwQ, and DeepSeek-R1 in prima.cpp. The results are given in Table 6.
Figure 7: Comparison of token latency and TTFT for llama.cpp, exo, dllama, and prima.cpp.
Table 6: Token latency (in millisecond/token) for Qwen, QwQ and distilled DeepSeek R1 models.
Model llama.cpp exo1 dllama prima.cpp
Qwen-2.5-7B 14 86 - 44
DeepSeek-R1-Distill-Qwen-7B 14 682 - 52
DeepSeek-R1-Distill-Llama-8B 14 772 435 59
Qwen-2.5-14B 23 317103 - 65
DeepSeek-R1-Distill-Qwen-14B 24 234753 - 76
@@
    Exo supports only the MLX backend for these models, so only D1 works.
  • 2
@@
    Latency decreases because exo provides full-precision Qwen models but 3-bit quantized DeepSeek models.
  • 3
    Latency spikes because exo is swapping data in/out from disk.
A.5 Select Devices to Build the Most Powerful Cluster
Existing systems require the cluster with enough RAM/VRAM, users have to gather more devices to run larger models. However, collecting enough devices is challenging. Here we raise two questions: (a) should we collect enough devices to meet the model’s needs? (b) Is more devices always better? The answer is no. Figure 8 shows how layer assignment adjusts and token latency varies on Llama 3-70B as devices reduce from 6 to 1.
Figure 8: Layer assignment and token latency over different number of devices.
@@ A.5 Select Devices to Build the Most Powerful Cluster
For question (a), with only 3 devices (D2, D3, and D5), the total available RAM+VRAM is 38 GiB, which is insufficient to hold the 40 GiB Llama 3-70B (Q4K) model. However, thanks to the fast SSDs on D2 and D3, mmap can swap model layers in a flash, and prima.cpp achieves the lowest latency. Thus, prima.cpp does not require enough memory to hold the entire model.
@@
For question (b), when we increase the number of devices to 6, the total available RAM+VRAM reaches 50 GiB, enough to hold the entire model. However, token latency is lower with just 3 devices, as the additional devices (D4 and D6) have weak CPUs and slow disks, creating bottlenecks. This shows that more devices do not always result in faster inference.
@@
This raises a new question: (c) If user has a device with a weak CPU or a slow disk, should it be added to the cluster? Intuitively, such a weak device would be a bottleneck. However, in cases of severe memory shortage, it actually help. For example, D2 and D3 are devices with a GPU, strong CPU, and fast disk, while D5 is a weak device. As shown in Figure 8, adding D5 reduced token latency by half, as the disk loading latency on D3 (which was heavily overloaded) became more problematic than D5’s computing. This raises more complex questions: if users have some weak devices, which ones should be added? More generally, (d) given a set of heterogeneous devices, how can we select a subset to build the best-performing cluster? This is challenging due to the uncertain number of devices to be selected, the highly heterogeneous cluster, and the various factors like CPU, GPU, RAM, VRAM, disk, network, and even OS that significantly affect inference speed. Fortunately, prima.cpp offers an easy solution: start by including all devices in the cluster, then remove those with only one assigned layer or fewer than a set threshold, as Halda identifies them as drags. Future updates will automate this process for easier to use.
A.6 Efficient Workload Distribution and Memory Usage
@@
Figure 9 shows each device’s RAM and VRAM usage to illustrate why prima.cpp achieves faster speed and prevents OOM. As exo and dllama don’t support Llama 14B-65B and encounter OOM at 70B, the memory usage for 14B-70B in Figures 9(b) and 9(c) is estimated based on system behavior and memory load at 8B. Figure 9(a) has been discussed in Section 4.1. Figures 9(b) and 9(c) show that exo and dllama consume high memory. Exo mixes multiple backends, using MLX on macOS for 4-bit computation and Tinygrad on Linux, where model weights load in 16-bit on the CPU and decoded to 32-bit on the GPU. In our case, D1(8 GiB UMA RAM) and D2 (8 GiB VRAM) get the same number of model layers, yet D2-CPU uses 4× more RAM and D2-GPU 8× more VRAM than D1-GPU. This results in high memory usage on Linux devices, increasing the risk of OOM. For dllama, it uses tensor parallelism and Q40 quantization to distribute and compress memory usage but lacks GPU support, so all memory load is on RAM and inference speed is limited. It has similar memory usage across devices due to its uniform tensor splitting, which causes problems on low-memory devices. In Figure 9(c), when running a 30B model, D2/D3 have more available RAM, while D1/D4 have less. To allocate enough memory, D1/D4 must free more active pages or swap out app data, slowing user apps or even the system. In such cases, OOM might be the better outcome. Additionally, D3 (the head device) loads the entire model before slicing and distributing it, taking significant RAM and making it more prone to OOM.
@@
In contrast, prima.cpp optimizes workload distribution with Halda and prevents memory waste with mmap. Though the solution to (1-5) is unobvious, we can observe Halda’s preference from Figure 9(d): powerful GPUs $>$ weak GPUs $>$ powerful CPUs $>$ fast disks. For example, at 8B–30B, Halda first fills D2-GPU and D3-GPU. At 45-65B, it fills D1-CPU to D4-CPU. Lastly, the remaining layers are placed on D2-CPU and D3-CPU because they have fast disks. This assignment prevents weak CPUs and slow disks from being used. Finally, only D2-CPU and D3-CPU experience RAM overload, but this does not cause OOM because the OS will free inactive mmap-ed pages instantly and prefetch model layers in advance. With fast disk reads, disk loading latency stays low, ensuring minimal token latency, which is exactly the result of our optimization goal (1). Beyond the advanced workload distribution, prima.cpp also prevents memory waste. With mmap, it loads only required model layers instead of the full model, eliminating the need for model slicing. Additionally, it supports model inference in Q4K format across heterogeneous platforms, eliminating the need to decode back to 16/32-bit, so RAM/VRAM usage is further reduced.
Figure 9: Comparison of memory usage on each device for llama.cpp, exo, dllama, and prima.cpp.

### v2

@@ A.3 Layer-to-Device Assignment: From Latency Analysis to Vectorized Model
To apply constraints to the subset of $\bm{w}$} and $\bm{n}$} corresponding to $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3},\mathcal{M}_{4}$, we define diagonal matrices ${\bm{P}}_{w}=\operatorname{diag}(-{\bm{I}}_{\mathcal{M}_{1}},-{\bm{I}}_{\mathcal{M}_{2}},-{\bm{I}}_{\mathcal{M}_{3}},{\bm{P}}_{\mathcal{M}_{4}}^{1},{\bm{P}}_{\mathcal{M}_{4}}^{2},{\bm{P}}_{\mathcal{M}_{4}}^{3})$, ${\bm{P}}_{n}=\operatorname{diag}(\mathbf{0}_{\mathcal{M}_{1}},\mathbf{0}_{\mathcal{M}_{2}},{\bm{I}}_{\mathcal{M}_{3}},\mathbf{0}_{\mathcal{M}_{4}}^{1},\mathbf{0}_{\mathcal{M}_{4}}^{2},-\mathbf{P}_{\mathcal{M}_{4}}^{3})$, where ${\bm{I}}_{\mathcal{M}_{1}},{\bm{I}}_{\mathcal{M}_{2}},{\bm{I}}_{\mathcal{M}_{3}}$ are identity matrices and $\mathbf{0}_{\mathcal{M}_{1}},\mathbf{0}_{\mathcal{M}_{2}},\mathbf{0}_{\mathcal{M}_{3}}$ are zero matrices corresponding to the subsets $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$, and ${\bm{P}}_{\mathcal{M}_{4}}^{1},{\bm{P}}_{\mathcal{M}_{4}}^{2},{\bm{P}}_{\mathcal{M}_{4}}^{3}$ are diagonal binary matricies (i.e., selection matricies) corresponding to the three constraints (30-32) within the subset $\mathcal{M}_{4}$. To construct ${\bm{P}}_{\mathcal{M}_{4}}^{1},{\bm{P}}_{\mathcal{M}_{4}}^{2},{\bm{P}}_{\mathcal{M}_{4}}^{3}$, we define a binary vector ${\bm{p}}_{\text{macOS}}$, where a value of 1 indicates that the current device is running on macOS and a value of 0 indicates otherwise. The number of elements in ${\bm{p}}_{\text{macOS}}$ matches the number of devices in the set $\mathcal{M}_{4}$. Similarly, we define binary vectors ${\bm{p}}_{\text{Linux}}$, ${\bm{p}}_{\text{Android}}$, ${\bm{p}}_{\text{metal}}$. Thus, we have ${\bm{P}}_{\mathcal{M}_{4}}^{1}=\operatorname{diag}({\bm{p}}_{\text{macOS}}\odot(1-{\bm{p}}_{\text{metal}}))$, ${\bm{P}}_{\mathcal{M}_{4}}^{2}=\operatorname{diag}({\bm{p}}_{\text{macOS}}\odot{\bm{p}}_{\text{metal}})$, ${\bm{P}}_{\mathcal{M}_{4}}^{3}={\bm{p}}_{\text{Linux}}+{\bm{p}}_{\text{Android}}$.
To handle constraints (34-35), we define ${\bm{P}}_{n}^{\text{gpu}}$ as a similar diagonal binary matrix, with elements set to one for devices with CUDA or Metal support. Specifically, we let ${\bm{P}}_{n}^{\text{gpu}}={\bm{P}}_{n}^{\text{cuda}}+{\bm{P}}_{n}^{\text{metal}}$, where ${\bm{P}}_{n}^{\text{cuda}}=\operatorname{diag}(\mathbf{0}_{\mathcal{M}_{1}},\mathbf{0}_{\mathcal{M}_{2}},{\bm{P}}_{\mathcal{M}_{3}}^{\text{cuda}},{\bm{P}}_{\mathcal{M}_{4}}^{\text{cuda}})$ and ${\bm{P}}_{n}^{\text{metal}}=\operatorname{diag}(\mathbf{0}_{\mathcal{M}_{1}},{\bm{I}}_{\mathcal{M}_{2}},\mathbf{0}_{\mathcal{M}_{3}},{\bm{P}}_{\mathcal{M}_{4}}^{\text{metal}})$.
@@
and the VRAM/shared-memory upper bound be ${\bm{z}}^{\text{gpu}}=[z_{1}^{\text{gpu}},\cdots,z_{M}^{\text{gpu}}]$, where
$z_{m}^{\text{gpu}}=\frac{1}{Lb^{\prime}}\cdot\begin{cases}0,&\text{if}\ \ \mathbb{I}_{\text{cuda}}=0\,\,\text{and}\,\,\mathbb{I}_{\text{metal}}=0,\\ d_{m,\text{cuda}}^{\text{avail}}-c^{\text{gpu}},&\text{if}\ \ \mathbb{I}_{\text{cuda}}=1,\\ d_{m,\text{metal}}^{\text{avail}}-c^{\text{gpu}},&\text{if}\ \ \mathbb{I}_{\text{metal}}=1\,\,\text{and}\,\,m\neq 1,\\ d_{m,\text{metal}}^{\text{avail}}-c^{\text{gpu}}-b_{o},&\text{if}\ \ \mathbb{I}_{\text{metal}}=1\,\,\text{and}\,\,m=1.\end{cases}$
The problem model can then be reformatted as:
@@
$b,b_{i},b_{o}$ Bytes of weight tensors per layer, and of input/output tensors.
$n_{kv}$ Number of tokens stored in the KV cache.
@@
$s_{\text{threshold}}^{\text{disk}}$ A threshold for disk read throughput. If the throughput is below this threshold, the disk is considered slow.
@@
${\bm{P}}_{w},{\bm{P}}_{n}$ Diagonal binary selection matrices for constraints on $\bm{w}$} and $\bm{n}$}.
${\bm{P}}_{n}^{\text{gpu}}$ Diagonal binary matrix that indicates whether a device uses a GPU.
@@
A.4 Generalization to a New Heterogeneous Testbed
To show the generalizability of prima.cpp, we repeated the experiments of Table 4 on another testbed. This testbed includes: a host PC (5 GB RAM, 1080 TI GPU with 11 GB VRAM), a Mac Mini (10 GB UMA RAM), a laptop (23 GB RAM, 3060 GPU with 6 GB VRAM), and a Redmi phone (7 GB RAM). Here, all RAM/VRAM values refer to available memory rather than total capacity. We ran llama.cpp on the laptop (with a 3060 GPU), and exo on the host PC, Mac Mini, and laptop (since exo is not supported on the phone). Meanwhile, dllama and prima.cpp ran on all four devices. Their TPOT values are listed in Table 7, following a trend similar to Table 4. In this case, llama.cpp encounters a VRAM bottleneck at 14B, and prima.cpp matches (only at 8B) or achieves the fastest speeds across 8-70B. Additionally, due to the limited RAM of the host PC, exo ran out of memory while loading the weight files. This supplementary experiment supports the generalizability of prima.cpp.
Table 7: TPOT (ms/token) on the new testbed.
Model llama.cpp exo dllama prima.cpp
Llama 3-8B 27 OOM 875 27
Llama 3-14B 199 - - 67
Llama 1-30B 469 - - 308
Llama 3-45B 623 - - 328
Llama 3-60B 12762 - - 671
Llama 1-65B 20073 - - 703
Llama 3-70B 23834 OOM OOM 718
Table 8: TPOT (ms/token) on a homogeneous testbed.
Model llama.cpp exo dllama prima.cpp
Llama 3-8B 15 OOM 81 15
Llama 3-14B 2243 - - 21
Llama 1-30B 6870 - - 52
Llama 3-45B 10563 - - 195
Llama 3-60B 14652 - - 391
Llama 1-65B 15798 - - 502
Llama 3-70B 17590 OOM OOM 1128
A.5 Generalization to a Homogeneous Testbed
We also compared llama.cpp, exo, dllama, and prima.cpp on a homogeneous, low-resource cluster. Since we do not have identical home devices, we used Docker to create four identical Linux containers on a server with an RTX 4090. Each container had 8 CPU cores, 8 GiB RAM (4 GiB available), 5 GiB VRAM, and 2 GB/s disk read throughput. Table 8 reports their TPOT results.
[…]
A.6 Run Prima.cpp on Llama 1&3, Qwen 2.5, QwQ and DeepSeek R1
Fig. 7 plots TPOT and TTFT for Llama models as curves. The prima.cpp curves (solid and dashed diamonds) consistently lie at the bottom, confirming that prima.cpp achieves the lowest TPOT and TTFT across 8-70B. Disabling prefetching or Halda increases latency, but the performance drop is substantially larger when Halda is disabled. This suggests that Halda contributes more critically to the speedup than prefetching.
Figure 7: TPOT and TTFT for Llama models across 8-70B.
Table 9: TPOT (ms/token) for Qwen 2.5, QwQ, and DeepSeek R1 models across 7-72B.
Model llama.cpp exo dllama prima.cpp
Qwen-2.5-7B 14 86 - 14
DeepSeek-R1-Distill-Qwen-7B 14 681 - 14
DeepSeek-R1-Distill-Llama-8B 14 771 435 14
Qwen-2.5-14B 23 31710 - 23
DeepSeek-R1-Distill-Qwen-14B 24 23475 - 24
@@ A.5 Generalization to a Homogeneous Testbed
    TPOT is lower because exo provides full-precision Qwen models, whereas the DeepSeek-distilled models are quantized to 3-bit.
[…]
A.7 Select Devices to Build the Most Powerful Cluster
Existing systems require clusters with sufficient aggregated RAM/VRAM, often forcing users to add more devices to support larger models. However, assembling enough devices is difficult for households. This raises two questions: (a) Should we collect enough devices to meet the model’s needs? (b) Do more devices always lead to better performance? The answer is no. Fig. 8 illustrates how the TPOT of prima.cpp on Llama 3-70B (Q4K) is affected by the device set. For each device set, the layer partitioning across device-backend pairs is determined by Halda.
Figure 8: Layer partitioning and TPOT over different device sets.
@@ A.7 Select Devices to Build the Most Powerful Cluster
For question (a), with only 3 devices (D2, D3, and D5), the aggregated RAM+VRAM is 38 GiB, which is insufficient to hold the 40 GiB Llama 3-70B (Q4K) model. However, thanks to the fast SSDs on D2 and D3, mmap can reload model layers quickly, and prima.cpp achieves the lowest latency. Thus, prima.cpp does not require memory sufficient to hold the entire model.
@@
For question (b), when we increase the number of devices to 6, the aggregated RAM+VRAM reaches 50 GiB, which is enough to hold the entire model. However, the TPOT is lower with just 3 devices, because the devices D4 and D6 have weak CPUs and slow disks, creating bottlenecks. This shows that more devices do not always result in faster inference.
This raises a new question: (c) If the user has a device with a weak CPU or a slow disk, should it be removed from the cluster? Intuitively, such a weak device would be a bottleneck. However, in cases of severe memory shortage, it does help. For example, D2 and D3 are devices with a GPU, a strong CPU, and a fast disk, while D5 is a weak device. As shown in Fig. 8, adding D5 reduced TPOT by roughly half because D3’s disk loading latency (which was heavily overloaded) dominated the impact of D5’s weaker CPU.
@@
This raises more questions: if users have some weak devices, which ones should be removed? More generally: (d) Given a set of heterogeneous devices, how can we select a subset to build the best-performing cluster? This is challenging due to the uncertain number of devices to be selected, the highly heterogeneous cluster, and the various factors like CPU, GPU, RAM, VRAM, disk, network, and even OS that significantly affect inference speed. Fortunately, Halda offers an easy solution: include all available devices, then remove those with only one assigned layer, since Halda indicates that excluding them would improve speed. This procedure is made automatically in prima.cpp.
A.8 Memory Footprint on Device-backend Pairs
Fig. 9 shows each device’s RAM and VRAM usage to illustrate why prima.cpp achieves faster speed and prevents OOM. As exo and dllama don’t support Llama 14B-65B and encounter OOM at 70B, the memory usage for 14B-70B in Figs. 9(b) and 9(c) is estimated based on system behavior and memory load at 8B. Fig. 9(a) was discussed in Section 4.1.
[…]
@@ A.8 Memory Footprint on Device-backend Pairs
For dllama, it uses TP and Q40 quantization to distribute and compress memory usage, but lacks GPU support, so all memory load is on RAM, and inference speed is limited. It has similar memory usage across devices due to its uniform tensor splitting, which causes problems on low-memory devices. In Fig. 9(c), for a 30B model, D2 and D3 have more available RAM, while D1 and D4 have less. To allocate enough memory, D1 and D4 must free more active pages or swap out application data, which can slow user apps or even crash the system. In such cases, OOM may be the safer outcome. Additionally, D3 (the head device) loads the entire model before slicing and distributing it, taking significant RAM and making it more prone to OOM.
@@
In contrast, prima.cpp optimizes workload distribution with Halda and prevents OOMs with mmap. Although the solution to the problem Eqs. (1)-(5) is hard to understand, we can observe Halda’s preference from Fig. 9(d): powerful GPUs > weak GPUs > powerful CPUs > fast disks. For example, at 8B-30B, Halda first fills D2-GPU and D3-GPU. At 45-65B, it fills D1-CPU to D4-CPU. Lastly, the remaining layers are placed on D2-CPU and D3-CPU because they have fast disks. This assignment prevents weak CPUs and slow disks from being used. Finally, only D2-CPU and D3-CPU experience RAM overload, but this does not cause OOM because the OS will free inactive mmapped pages instantly and prefetch model layers in advance. With fast disk reads, disk loading latency stays low, ensuring minimal TPOT, which is exactly the result of our optimization goal (Eq. 1).
Beyond the advanced workload distribution via Halda, prima.cpp also prevents memory waste. With mmap, it loads only the required model layers instead of the full model, eliminating the need for model slicing (which may cause OOMs like in dllama). Additionally, it supports model inference in Q4K format across heterogeneous platforms, eliminating the need to decode back to 16-bit or 32-bit, thereby further reducing RAM/VRAM usage.
A.9 Run Prima.cpp with Speculative Decoding
Prima.cpp can be further accelerated with speculative decoding. In prima.cpp, we add support for this technique. Since the draft model is small (0.5-3B), we run it as a standalone process on the head device and set the most powerful device as the head. The draft model predicts 5 tokens per step, which are then verified in batch by the larger target model. The testbed consists of 4 Linux devices, each with an 8-core CPU, 8 GiB RAM, and a sequential disk read throughput of 600 MB/s. Two nodes are equipped with a 4090 GPU, but each is limited to 11 GiB VRAM.
Table 10 compares the TPOT of llama.cpp and prima.cpp with and without speculative decoding. With speculative decoding, prima.cpp delivers an additional 25-45% latency reduction across 14-70B models. For example, Qwen 2.5-32B speeds up from 18 to 26 tokens/s, and Llama 3.3-70B from 1.2 to 2.3 tokens/s. At this throughput, a 32B model meets the 20-50 tokens/s throughput commonly required by LLM agents, facilitating broader deployment of frontier LLM agents on home devices.
Figure 9: Memory footprint on each device-backend pair.
Table 10: TPOT (ms/token) for llama.cpp and prima.cpp with and without speculative decoding.
Model llama.cpp prima.cpp prima.cpp (with speculative)
Qwen-2.5-7B $20\pm 0$ ms $20\pm 0$ ms $18\pm 1$ ms (draft 0.5B on GPU)
Llama 3.2-8B $20\pm 0$ ms $20\pm 0$ ms $20\pm 1$ ms (draft 1B on GPU)
Qwen-2.5-14B $36\pm 1$ ms $36\pm 0$ ms $27\pm 1$ ms (draft 1.5B on GPU)
DeepSeek-R1-Distill-Qwen-14B $32\pm 1$ ms $32\pm 0$ ms $22\pm 1$ ms (draft 1.5B on GPU)
Qwen-2.5-32B $6551\pm 38$ ms $55\pm 0$ ms $38\pm 5$ ms (draft 0.5B on GPU)
DeepSeek-R1-Distill-Llama-70B $20118\pm 69$ ms $859\pm 40$ ms $593\pm 65$ ms (draft 8B on CPU)
Llama 3.3-70B $20083\pm 10$ ms $803\pm 9$ ms $442\pm 11$ ms (draft 3B on GPU)
Qwen-2.5-72B $21600\pm 80$ ms $963\pm 18$ ms $544\pm 19$ ms (draft 3B on CPU)
A.10 Ablation Study of Halda with Heuristic Scheduler Baselines
While prior work has studied layer partitioning across heterogeneous devices, these approaches often rely on strong and unrealistic assumptions that limit their use in real home clusters: (a) They only support workload scheduling within GPU clusters; (b) They require the cluster’s aggregated VRAM to meet the model’s needs; (c) They assume all devices are necessary. For (a) and (b), most households cannot afford an expensive machine with sufficient VRAM, let alone a GPU cluster. For (c), since home devices can vary widely, it could be better to drop weak devices than to keep them.
The novelty in Halda lies in supporting GPU/CPU-mixed clusters while relaxing memory requirements, and in explicitly accounting for disk latency and OS-specific memory behaviors. Moreover, Halda can automatically select the optimal subset of devices from a candidate pool to serve an inference engine that runs at maximum speed. We call this ability "device selection". No previous work has considered these features, but they are necessary in home clusters.
To better demonstrate Halda’s effectiveness and novelty, we conduct a supplementary experiment comparing it with two heuristic scheduler baselines:
  • MemSched: Partitions layers according to RAM/VRAM ratios. This method originates from exo and is also the default strategy in prima.cpp (w/o halda).
  • PerfSched: This method originates from Galaxy. First, partition layers by devices’ compute power, then migrate OOM layers to other devices. This can still hit OOM, so we add a greedy fallback: If OOM persists, offload to CPUs (in proportion to CPU compute power).
The testbed uses a Mac Mini (16GB RAM), a host PC (16GB RAM and 11GB 1080TI GPU), a laptop (32GB RAM and 6GB 3060 GPU), and a Redmi phone (16GB RAM). In this setup, aggregated VRAM is insufficient, but aggregated RAM+VRAM is sufficient, so OOM does not occur.
Figure 10: TPOT of MemSched, PerfSched, and Halda on Llama models across 8-70B.
[…]